# Yixin Liu1 2  Lie Lu2  Jihui Jin2  Lichao Sun1  Andrea Fanelli2

### v2

@@ 4 Methodology
Despite AudioSeal achieving robust detection, the model struggles to perform accurate message decoding even without distortion (See Table 5). In this work, following the architecture backbone of AudioSeal, we identify and resolve two architectural limitations: disjoint generator-detector and the information bottleneck caused by embedding mean-pooling. Specifically, we first propose a cross-attention generator-detector watermarking system with a shared embedding table and temporal modulation to improve learning efficiency in message decoding. Then, to further improve watermark quality, we propose a psychoacoustic-aligned temporal-frequency masking $\ell_{2}$ loss. The overall framework is shown in Figure 2.
@@ 4.1 Cross-Attention Generator-Detector Watermarking System with Shared Embedding Table
We observe that the fully disjointed architecture of AudioSeal ($\Theta_{\mathcal{G}}\neq\Theta_{\mathcal{D}}$) often converges fast for the watermark detection learning but struggles to learn the message decoding part efficiently and accurately. On the other hand, the full parameter-sharing architecture of WavMark ($\Theta_{\mathcal{G}}=\Theta_{\mathcal{D}}$) can achieve superior efficiency in learning both detection and message decoding but lacks enough robustness capability against various distortions (See App. C.2). This motivates us to explore a blended architecture with partial parameter sharing between the generator and the detector, which has the potential to boost message decoding learning while preserving robust capability.
@@
Our key design is to share the message conditioning module $\mathcal{M}$ between the generator and the detector, which helps bridge the information flow between how the message is composed and how it can be reconstructed in the detector part. Specifically, the learnable part of the embedding table $\mathbf{E}$ in $\mathcal{M}$, which serves as the fundamental vector set for composing the message latent in the generator part, is now utilized as a reference when decomposing the latent to retrieve the message bits in the detector. To achieve this, we propose leveraging a cross-attention block (Vaswani, 2017) to use the embedding table $\mathbf{E}$ as a (key-value) reference for message decoding, given the query audio latent $\bm{\tilde{h}}_{x}$. We describe this module in detail in the following section.
@@
Message Decoding with Cross-Attention. Ideally, for the given embedding table $\mathbf{E}$ with $2K$ entries and a watermarked audio $\bm{x}_{w}=\bm{x}+\mathcal{G}(\bm{x},\bm{w})$, we want to reconstruct the original $K$ embedding vectors $[\bm{E}_{I_{1}},\dots,\bm{E}_{I_{K}}]$, that were used to compose the ground-truth message latent $\bm{h}_{w}$ and then feed them as context for the message decoding. To achieve this, we use an attention mechanism that transforms the embedding table into key and value using two different linear projections, and does attention-based vector merging with a query from the reconstructed latent representation $\bm{\tilde{h}}_{x}$. Specifically, since the nearby two entries in $\mathbf{E}$ represent one position with different bit states, merging them into one $2H$-dim vector represents the latent of each position. We transform this reshaped embedding matrix $\mathbf{E}^{\prime}\in\mathbb{R}^{K\times 2H}$ into $\mathbf{K},\mathbf{V}$ with two linear projections $\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{2H\times H}$, that is $\mathbf{K}=\mathbf{E}^{\prime}\mathbf{W}_{K}$ and $\mathbf{V}=\mathbf{E}^{\prime}\mathbf{W}_{V}$. Then, we “demodulate” along the temporal axis the reconstructed waveform latent $\bm{\tilde{h}}_{x}\in\mathbb{R}^{T\times H}$, to obtain the first version of raw prediction $\mathbf{\tilde{V}}_{x}$ for the original $K$ components. To do this, we first use a linear projection $\mathbf{W}_{dem}\in\mathbb{R}^{T\times K}$ for obtaining the query sequence $\mathbf{\tilde{h}}_{x}^{dem}=\mathbf{W}_{dem}^{T}\mathbf{\tilde{h}}_{x}\in\mathbb{R}^{K\times H}$ and follow with a linear query projection $\mathbf{W}_{Q}\in\mathbb{R}^{H\times H}$. Then we can utilize the embedding table $\mathbf{E}$ to further refine the final component prediction $\mathbf{\tilde{V}}_{x}$ with the following cross-attention mechanism:
$\mathbf{Q}$ $=\mathbf{\tilde{h}}_{x}^{dem}\,\mathbf{W}_{Q}\in\mathbb{R}^{K\times H},$ (2)
$\mathbf{K}$ $=\mathbf{E}^{\prime}\,\mathbf{W}_{K}\in\mathbb{R}^{K\times H},$
$\mathbf{V}$ $=\mathbf{E}^{\prime}\,\mathbf{W}_{V}\in\mathbb{R}^{K\times H},$
@@
$\mathbf{A}$ $=\mathrm{softmax}\!\Bigl{(}\frac{\mathbf{Q}\,\mathbf{K}^{\top}}{\sqrt{H}}\Bigr{)}\in\mathbb{R}^{K\times K},$
$\mathbf{\tilde{V}}_{x}$ $=\mathrm{act}\bigl{(}\mathbf{A}\,\mathbf{V}\bigr{)}\in\mathbb{R}^{K\times H},$
@@
where $H$ is the latent dimension, $K$ is the number of message bits, $t$ is the temporal dimension, and $\mathrm{act}(\cdot)$ can be any activation function. In this study, we use exponential linear units (ELU) Clevert (2015). With the predicted $\mathbf{\tilde{V}}_{x}$, we further employ a linear projection layer $\mathbf{W}_{dec}\in\mathbb{R}^{H\times 1}$ with sigmoid activation function for the construction of the message decoding head. The final logit for message prediction bit is then obtained as $\hat{\bm{w}}_{x}=\sigma\bigl{(}\mathbf{\tilde{V}}_{x}\mathbf{W}_{dec}\bigr{)}$, where $\sigma(\cdot)$ is the sigmoid activation function.
Message Embedding via Temporal Modulation. In AudioSeal, the message latent is obtained by a mean-pooling operation with temporal-axis repetition, $\bm{h}_{w}=\texttt{Repeat}\!\left(\frac{1}{K}\sum_{j=1}^{K}\bm{E}_{I_{j}},t^{\prime}\right)\in\mathbb{R}^{t^{\prime}\times H}.$
@@
This approach can be viewed as injecting the message information mostly into the frequency domain, which greatly limits the message hiding capabilities. To improve upon this, knowing the input audio length $T$, we introduce a temporal message conditioning mechanism that employs a linear modulation layer $\mathbf{W}_{M}\in\mathbb{R}^{K\times t^{\prime}}$ to obtain the message latent:
$\bm{h}_{w}=\mathbf{W}_{M}^{\top}\,\bm{V}(\bm{w})\;\in\;\mathbb{R}^{t^{\prime}\times H}.$ (3)
@@
This simple design not only prevents the sole reliance on the frequency domain for distributing the $K$-bit message, but also significantly facilitates the learning process for message decoding (see Figure 4).
4.2 Psychoacoustic-Aligned Temporal–Frequency Masking Loss
Achieving imperceptibility is a key requirement of any watermarking system. Ideally, a watermark should exploit the perception characteristics of the human auditory system so that artifacts remain imperceptible. Following AudioSeal, we place a $\ell_{1}$ constraint on the watermark signal $\delta_{x}$ to ensure waveform-domain smoothness, and we incorporate a multi-scale Mel spectrogram loss $\mathcal{L}_{\text{msspec}}$ (e.g., as in Defossez et al. (2022)) to manage frequency-domain fidelity.
We also adopt adversarial losses $\mathcal{L}_{adv}$ on multi-scale STFT spectrograms for perceptual improvement. However, simply pushing the watermarked audio distribution to be close to the clean audio in an adversarial sense may penalize desirable watermark characteristics (e.g., certain musical “remixes” can produce satisfactory imperceptible watermarks, but still yield high adversarial loss). Thus, we choose to down-weight $\mathcal{L}_{adv}$ and focus more on the psychoacoustic aspects to develop a human-centric, imperceptible watermark.
Psychoacoustic Masking Principles. Our design is based on psychoacoustic masking principles (Gelfand, 2017; Holdsworth et al., 1988), which describe how listeners are less sensitive to low-level sounds occurring near stronger sounds, both in time (forward/backward masking) and in frequency (simultaneous masking). In other words, when a loud audio component is present at a certain time–frequency location, it can mask weaker signals in its temporal and frequency vicinity (Necciari et al., 2016). Inspired by this, we propose a novel $\ell_{2}$ loss with per-tile TF-penalty.
@@ 4.2 Psychoacoustic-Aligned Temporal–Frequency Masking Loss
Each index pair $(m_{c},t_{c})\in\mathcal{M}_{\text{masker}}$ acts as a masker, and we model its masking energy over the neighboring tiles as a linear decay in the decibel domain. Specifically, given mel-scale frequency radius $r_{l}^{m}$, $r_{h}^{m}$, and time-axis radius $r_{b}^{t}$, $r_{f}^{t}$, the local region of a masker is defined as:
$R(m_{c},t_{c})=\{(m,t)\mid-r_{l}^{m}\leq\Delta m\leq r_{h}^{m},-r^{t}_{b}\leq\Delta t\leq r^{t}_{f}\},$
where $\Delta m=m-m_{c}$, and $\Delta t=t-t_{c}$. Aligned with the empirical studies (Necciari et al., 2016) that the post-masking region (forward masking) is usually longer than the pre-masking region (backward masking) in the temporal axis, we set asymmetric radii $r_{b}^{t}$ and $r_{f}^{t}$, with forward masking region of 200ms and backward masking region of 20ms. For the frequency-bin radius, we first compute the empirical critical bandwidth for each mel-bin using the formula from Zwicker & Fastl (2013), $W(m_{c})=25+75\left(1+1.4\left(\frac{\mathcal{F}(V[m_{c}])}{1000}\right)^{2}\right)^{0.69}$, where $\mathcal{F}(m_{c})$ is the approximated center frequency obtained by converting each Mel bin value $V[m_{c}]$ back to Hz using $\mathcal{F}(m_{c})=700\cdot(10^{\frac{V[m_{c}]}{2595}}-1)$. Then the radii for each mel-bin is set as $r_{l}^{m}=r_{h}^{m}=r_{b}^{m}\cdot\gamma$, where $\gamma=W(m_{c})/|\mathcal{F}(m_{1})-\mathcal{F}(m_{0})|$ is an adaptive scaling term, $V[m_{i}]$ denote the $i$-th mel-bin value, $r_{b}^{m}$ is a base radius that is set to be 3. With the dynamic radius scaled by the critical bandwidth – where a higher frequency component has a wider masking range (Holdsworth et al., 1988) – we observe a better performance than using constant $\gamma=1$.
The physical structure of the cochlea determines that lower frequencies (located at the base of the cochlea) cause broader and stronger excitation patterns compared to the higher frequencies (located at the apex). To inject this bias of slower decay in the upward spread of masking (Zwicker & Fastl, 2013), we set the decaying term for the upward direction to be larger than the downward direction, i.e., $\alpha_{f}^{+}\leq\alpha_{f}^{-}$. For the temporal decay, following Necciari et al. (2016), we set the backward direction to be larger than the forward direction, i.e., $\beta_{t}^{+}\leq\beta_{t}^{-}$. Given the masker $(m_{c},t_{c})$, the threshold energy $E_{\text{mask}}(m,t)$ for the maskee tile $(m,t)$ is computed as:
$E_{\text{mask}}(m,t;(m_{c},t_{c}))=20\log_{10}\mathcal{S}_{O}(m_{c},t_{c})-\Delta E,$ (4)
@@
This ensures each maskee tile receives masking energy from its most perceptually dominant masker, while weaker maskers have negligible impact (Gelfand, 2017).
Per-Tile-Weighted $\ell_{2}$ Loss Computation. Psychoacoustic masking effectively applies only when the masker’s masking energy surpasses the tile’s energy. We therefore apply the masking threshold on watermarked mel-spectrogram at $20\log_{10}\mathcal{S}_{W}(m,t)$, filtering the masked tiles set as:
@@
$\lambda(m,t)$ $=1+\mathbf{1}_{(m,t)\in\mathcal{M}_{\text{maskee}}}\cdot 10^{E^{*}_{\text{mask}}(m,t)/20},$ (5)

### v3

@@ 4 Methodology
Despite AudioSeal achieving robust detection, the model struggles to perform accurate message decoding even without distortion (See Table 7). In this work, following the architecture backbone of AudioSeal, we identify and resolve two architectural limitations: disjoint generator-detector and the information bottleneck caused by embedding mean-pooling. Specifically, we first propose a cross-attention generator-detector watermarking system with a shared embedding table and temporal modulation to improve learning efficiency in message decoding. Then, to further improve watermark quality, we propose a psychoacoustic-aligned TF masking $\ell_{2}$ loss. The framework is shown in Figure 2.
@@ 4.1 Cross-Attention Generator-Detector Watermarking System with Shared Embedding Table
We observe that the fully disjoint architecture of AudioSeal ($\Theta_{\mathcal{G}}\neq\Theta_{\mathcal{D}}$) often converges fast for the watermark detection learning but struggles to learn the message decoding part efficiently and accurately. On the other hand, the full parameter-sharing architecture of WavMark ($\Theta_{\mathcal{G}}=\Theta_{\mathcal{D}}$) can achieve superior efficiency in learning both detection and message decoding but lacks enough robustness capability against various distortions (See App. C.2). This motivates the design of a hybrid architecture with partial parameter sharing between the generator and the detector, which facilitates more effective message decoding while preserving robust capability.
@@
Our key design is to share the message conditioning module $\mathcal{M}$ between the generator and the detector, which helps bridge the information flow between how the message is composed and how it can be reconstructed in the detector part. Specifically, the learnable part of the embedding table $\bm{E}$ in $\mathcal{M}$, which serves as the fundamental vector set for composing the message latent in the generator part, is now utilized as a reference when decomposing the latent to retrieve the message bits in the detector. To achieve this, we propose leveraging a cross-attention block (Vaswani et al., 2017) to use the embedding table $\bm{E}$ as a (key-value) reference for message decoding, given the query audio latent $\bm{\tilde{h}}_{x}$. We describe the decoding mechanism in detail in the following section.
@@
Message Decoding with Cross-Attention. Ideally, for the given embedding table $\bm{E}$ with $2K$ entries and a watermarked audio $\bm{x}_{w}=\bm{x}+\mathcal{G}(\bm{x},\bm{w})$, we want to reconstruct the original $K$ embedding vectors $\bm{E}_{I_{1}},\dots,\bm{E}_{I_{K}}]$, that were used to compose the ground-truth message latent $\bm{h}_{w}$ and then feed them as context for the message decoding. To achieve this, we use an attention mechanism that transforms the embedding table into key and value using two different linear projections, and does attention-based vector merging with a query from the reconstructed latent representation $\bm{\tilde{h}}_{x}$. Specifically, since the nearby two entries in $\bm{E}$ represent one position with different bit states, merging them into one $2H$-dim vector represents the latent of each position. We transform this reshaped embedding matrix $\bm{E}^{\prime}\in\mathbb{R}^{K\times 2H}$ into $\bm{K},\bm{V}$ with two linear projections $\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{2H\times H}$, that is $\bm{K}=\bm{E}^{\prime}\bm{W}_{K}$ and $\bm{V}=\bm{E}^{\prime}\bm{W}_{V}$. Then, we demodulate the reconstructed waveform latent along the temporal axis $\bm{\tilde{h}}_{x}\in\mathbb{R}^{T\times H}$, to obtain the first version of raw prediction $\bm{\tilde{V}}_{x}$ for the original $K$ components. To do this, we first use a linear projection $\bm{W}_{dem}\in\mathbb{R}^{T\times K}$ for obtaining the query sequence $\bm{\tilde{h}}_{x}^{dem}=\bm{W}_{dem}^{T}\bm{\tilde{h}}_{x}\in\mathbb{R}^{K\times H}$ and follow with a linear query projection $\bm{W}_{Q}\in\mathbb{R}^{H\times H}$. Then we can utilize the embedding table $\bm{E}$ to further refine the final component prediction $\bm{\tilde{V}}_{x}$ with the following cross-attention mechanism:
$\bm{Q}$ $=\bm{\tilde{h}}_{x}^{dem}\,\bm{W}_{Q}\in\mathbb{R}^{K\times H},$ (2)
$\bm{K}$ $=\bm{E}^{\prime}\,\bm{W}_{K}\in\mathbb{R}^{K\times H},$
$\bm{V}$ $=\bm{E}^{\prime}\,\bm{W}_{V}\in\mathbb{R}^{K\times H},$
@@
$\bm{A}$ $=\mathrm{softmax}\!\Bigl(\frac{\bm{Q}\,\bm{K}^{\top}}{\sqrt{H}}\Bigr)\in\mathbb{R}^{K\times K},$
$\bm{\tilde{V}}_{x}$ $=\mathrm{act}\bigl(\bm{A}\,\bm{V}\bigr)\in\mathbb{R}^{K\times H},$
@@
where $H$ is the latent dimension, $K$ is the number of message bits, $t$ is the temporal dimension, and $\mathrm{act}(\cdot)$ can be any activation function. In this study, we use exponential linear units (ELU) (Clevert et al., 2016). With the predicted $\bm{\tilde{V}}_{x}$, we further employ a linear projection layer $\bm{W}_{dec}\in\mathbb{R}^{H\times 1}$ with sigmoid activation function for the construction of the message decoding head. The final logit for message prediction bit is then obtained as $\hat{\bm{w}}_{x}=\sigma\bigl(\bm{\tilde{V}}_{x}\bm{W}_{dec}\bigr)$, where $\sigma(\cdot)$ is the sigmoid activation function.
@@
Message Embedding via Temporal Modulation. In AudioSeal, the message latent is obtained by a mean-pooling operation with temporal-axis repetition, $\bm{h}_{w}=\texttt{Repeat}\!\left(\frac{1}{K}\sum_{j=1}^{K}\bm{E}_{I_{j}},t^{\prime}\right)\in\mathbb{R}^{t^{\prime}\times H}.$ This approach can be viewed as injecting the message information mostly into the frequency domain, which greatly limits the message hiding capabilities. To improve upon this, knowing the input audio length $T$, we introduce a temporal message conditioning mechanism that employs a linear modulation layer $\bm{W}_{M}\in\mathbb{R}^{K\times t^{\prime}}$ to obtain the message latent:
$\bm{h}_{w}=\bm{W}_{M}^{\top}\,\bm{V}(\bm{w})\;\in\;\mathbb{R}^{t^{\prime}\times H}.$ (3)
@@
This design avoids relying only on the frequency domain for distributing the $K$-bit message, but also significantly facilitates the learning process for message decoding (see Figure 6).
4.2 Psychoacoustic-Aligned Time-Frequency Masking
Achieving imperceptibility is a key requirement of any watermarking system. Ideally, a watermark should exploit the perception characteristics of the human auditory system so that the artifacts remain imperceptible. Among all the psychoacoustic effects, auditory masking (Gelfand, 2017; Holdsworth et al., 1988) offers strong theoretical guidance for designing perceptual losses in watermarking (Deng et al., 2023; Qin et al., 2019). While prior work like AudioSeal (San Roman et al., 2024) introduced perceptual losses such as TF-Loudness, these often adopt a coarse approach based on loudness differences within fixed TF tiles. Such methods face two main limitations: (i) they may not fully capture sophisticated auditory masking effects, particularly the interactions between masker and maskee signals across different tiles, and (ii) loudness differences alone can provide weak supervisory signals for guiding the watermark embedding, resulting in a coarse penalty distribution (visualized in App. C.9 in Figure 12). These limitations motivate our development of a more psychoacoustic-aligned TF-masking loss.
@@ 4.2 Psychoacoustic-Aligned Time-Frequency Masking
Each index pair $(m_{c},t_{c})\in\mathcal{M}_{\text{masker}}$ acts as a masker, and we model its masking influence over neighboring tiles with a linear energy decay in the decibel domain. Specifically, given mel-scale frequency radii $r_{l}^{m}$, $r_{h}^{m}$, and time-axis radius $r_{b}^{t}$, $r_{f}^{t}$, the local region of a masker is defined as:
$R(m_{c},t_{c})=\{(m,t)\mid-r_{l}^{m}\leq\Delta m\leq r_{h}^{m},-r_{b}^{t}\leq\Delta t\leq r_{f}^{t}\},$
where $\Delta m=m-m_{c}$, and $\Delta t=t-t_{c}$. Aligned with the empirical studies (Necciari et al., 2016) that the post-masking region (forward masking) is usually longer than the pre-masking region (backward masking) in the temporal axis, we set asymmetric radii $r_{b}^{t}$ and $r_{f}^{t}$, with a forward masking region of 200ms and a backward masking region of 20ms. For the frequency-bin radius, we first compute the empirical critical bandwidth for each mel-bin using the formula from Zwicker and Fastl (2013), $W(m_{c})=25+75\left(1+1.4\left(\frac{\mathcal{F}(m_{c})}{1000}\right)^{2}\right)^{0.69}$, where $\mathcal{F}(m_{c})$ is the approximated center frequency obtained by converting each Mel bin value $m_{c}$ back to Hz using $\mathcal{F}(m_{c})=700\cdot(10^{\frac{m_{c}}{2595}}-1)$. Then the radii for each mel-bin are set as $r_{l}^{m}=r_{h}^{m}=r_{b}^{m}\cdot\gamma$, where $\gamma=W(m_{c})/|\mathcal{F}(m_{1})-\mathcal{F}(m_{0})|$ is an adaptive scaling term, $V[m_{i}]$ denote the $i$-th mel-bin value, $r_{b}^{m}$ is a base radius that is set to be 3. With the dynamic radius scaled by the critical bandwidth – where higher‑frequency components have wider masking ranges (Holdsworth et al., 1988) – we observe a better performance than using constant $\gamma$.
The physical structure of the cochlea determines that higher frequencies (located at the base of the cochlea) cause broader and stronger excitation patterns compared to the lower frequencies (located at the apex). To inject this bias of slower decay in the upward spread of masking (Zwicker and Fastl, 2013), we set the decaying term for the upward direction to be smaller than the one in the downward direction, i.e., $\alpha_{f}^{+}\leq\alpha_{f}^{-}$. For the temporal decay, following Necciari et al. (2016), since the forward decay is slower, we set the forward decaying slope less steep than the backward one, i.e., $\beta_{t}^{+}\leq\beta_{t}^{-}$. Given a masker $(m_{c},t_{c})$, the threshold energy $E_{\text{mask}}$ for the maskee tile $(m,t)$ is computed as:
[…]
@@
This ensures that each maskee tile receives masking energy from its most perceptually dominant masker, while weaker maskers have negligible impact (Gelfand, 2017).
Per-Tile-Weighted $\ell_{2}$ Loss Computation. Psychoacoustic masking effectively applies only when the masker’s masking energy surpasses the tile’s energy. We therefore apply the masking threshold on watermarked mel-spectrogram at $20\log_{10}\mathcal{S}_{W}(m,t)$, filtering the set of masked tiles as:
@@
$\lambda(m,t)$ $=1+\bm{1}_{(m,t)\in\mathcal{M}_{\text{maskee}}}\cdot 10^{E^{*}_{\text{mask}}(m,t)/20},$ (5)
@@
Other Auxiliary Perceptual Loss. Following San Roman et al. (2024), we use a $\ell_{1}$ constraint on the watermark signal $\delta_{w}$ to ensure waveform-domain smoothness, and we incorporate a multi-scale Mel spectrogram loss $\mathcal{L}_{\text{msspec}}$ (e.g., as in Défossez et al. (2023)) to manage frequency-domain fidelity. Furthermore, we also adopt adversarial losses $\mathcal{L}_{adv}$ on multi-scale STFT spectrograms for perceptual improvement, but with a lower weight $\lambda_{\text{adv}}=1$ (while AudioSeal uses $\lambda_{\text{adv}}=4$). The reason behind this is that aggressively pushing the watermarked audio distribution to be close to the clean audio distribution in an adversarial sense may penalize desirable watermark characteristics (e.g., certain musical “remixes” can produce satisfactory imperceptible watermarks, but still yield high adversarial loss). We observe that lowering the adversarial weight allows the model to focus more on the psychoacoustic aspects to develop a human-centric, imperceptible watermark.